# From Counterfactuals to Trees: Competitive Analysis of Model Extraction Attacks

### v2

@@ 1 Introduction
While MLaaS platforms provide accessible and scalable ML solutions, they must address the growing demand for explainability in their decision-making processes. Regulatory frameworks such as the EU AI Act’s Article 13111https://artificialintelligenceact.eu/article/13/ further mandate greater transparency across a wide range of applications. In response, MLaaS providers increasingly incorporate explainability techniques to elucidate model behavior and ensure fairness. However, studies have shown that querying a model’s explanations can enable attackers to replicate its parameters and architecture, effectively copying the original model [Tramèr et al., 2016, Wang et al., 2022, Aïvodji et al., 2020, Öksüz et al., 2024]. This reveals a critical tension between the need for transparency and the protection of model integrity and intellectual property.
@@
Robot explorationRefer to caption“Line of sight”Model $\bullet$×$\times$דCounterfactualquery”
@@ 2 Online Discovery, Model Extraction Attacks and Competitive Analysis
Competitive Analysis. A central tool for analyzing online discovery problems is competitive analysis [Karlin et al., 1986], which compares the performance of an online algorithm — one that adapts its decisions based solely on information acquired so far — to an optimal offline algorithm with complete foresight. Formally, we measure the ratio between: (i) the complexity (e.g., number of queries, computational cost) incurred by the online algorithm to achieve its goal and (ii) the minimal complexity that an offline algorithm, with complete foresight, would require to accomplish the same task or verify its reconstruction. A constant ratio implies a constant-competitive algorithm; in more complex settings, the ratio may grow with problem parameters.
@@ 3.2 Tree Reconstruction Attack algorithm
We now introduce our proposed Tree Reconstruction Attack (TRA), detailed in Algorithm 1. TRA is a divide-and-conquer based algorithm that aims to reconstruct a decision tree $f_{n}$ with $n$ split levels by systematically exploring the input space $\mathcal{X}$, using only a black-box API returning predictions and locally optimal counterfactuals. A split level is defined as a particular value for a given feature that divides the input space into two subspaces. Note that multiple nodes within different branches of a decision tree can share the same split level.
Algorithm Overview. TRA operates by maintaining a query list $\mathcal{Q}$ that initially contains the entire input space $\mathcal{X}$. The algorithm iteratively processes each input subset $\mathcal{E}\subseteq\mathcal{X}$ from $\mathcal{Q}$, until $\mathcal{Q}$ is empty, ensuring that all decision boundaries of the target model $f$ are identified and replicated in the reconstructed tree. More precisely, at each iteration, TRA first retrieves the subset $\mathcal{E}$ on top of the priority queue $\mathcal{Q}$. It computes its geometric center $x$ using the $center$ function (line 5). It then queries the oracle $\mathcal{O}_{d}$ with the target model $f$, input $x$, and subset $\mathcal{E}$ to obtain a counterfactual explanation $x^{\prime}=\mathcal{O}_{d}(f,x,\mathcal{E})$ (line 8). The set of feature indices where $x^{\prime}$ differs from $x$, i.e., $\{i\mid x^{\prime}_{i}\neq x_{i}\}$ is consequently identified. For each differing feature $i$, TRA splits the input subset $\mathcal{E}$ into two subspaces based on the split value $x^{\prime}_{i}$ ($split$ function). The resulting subspaces are added to $\mathcal{Q}$ for further exploration (line 9). If no counterfactual explanation $x^{\prime}$ exists within $\mathcal{E}$ , TRA assigns the label $y=f(x)$ to $\mathcal{E}$, indicating that it corresponds to a leaf node in the reconstructed tree (line 11).
@@
2:  $\mathcal{Q}\leftarrow\{\mathcal{X}\}$Initialize query list with the entire input space}
@@
9:        $\mathcal{Q}\leftarrow insert(split(\mathcal{E},x,x^{\prime}))$Split $\mathcal{E}$ based on $x^{\prime}$ and add the resulting subspaces to $\mathcal{Q}$}
@@
Let $f_{n}$ be a decision tree with $n$ split levels across a $m$-dimensional input space $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{m}$. Denote $s_{i}$ as the number of split levels in $f_{n}$ over the $i$-th feature, such that $\sum_{i=1}^{m}s_{i}=n$. The worst-case complexity of Algorithm 1 is $O\left(\prod_{\begin{subarray}{c}i=1\end{subarray}}^{m}(s_{i}+1)\right)$.
@@
The worst-case complexity of Algorithm 1 is $O\left(\left(1+\frac{n}{m}\right)^{m}\right)$.
The proofs of Proposition 3.4 and Corollary 3.5 are provided in the Appendix A. Proposition 3.4 establishes a first simple upper bound on the complexity of Algorithm 1. Intuitively, consider a two-dimensional decision boundary that partitions the space in a chessboard-like pattern of size $s_{1}\times s_{2}$. A comprehensive mapping of such space necessitates at least $s_{1}\times s_{2}$ queries (one in each sub-square), a requirement that holds for multi-class classification scenarios and in high-dimension.
@@
$C_{TRA}^{(n,m)}=\frac{\sum_{i=1}^{m}s_{i}\prod_{j=1}^{i-1}(s_{j}+1)+\prod_{j=1}^{m}(s_{j}+1)}{2n+1}\leq\frac{2\left(1+\frac{n}{m}\right)^{m}-1}{2n+1},$
@@
The proofs of Propositions 3.6 and 3.7 are provided in Appendix A. Proposition 3.6 provides the competitive ratio achieved by the TRA algorithm, while Proposition 3.7 establishes that the choice of query position does not affect the competitive ratio for any divide-and-conquer algorithm iteratively partitioning the input space. These propositions demonstrate that TRA not only offers a competitive approach to tree reconstruction under various query selection strategies, but also sets a theoretical benchmark that cannot be surpassed by other divide-and-conquer-based methods.
Anytime Behavior. Since the query budget is often not known in advance and may vary depending on the target model, it is crucial for an extraction attack to operate in an anytime fashion—that is, to produce a usable classifier even if interrupted before completion. TRA satisfies this property by assigning provisional labels at each input space split (line 9 of Algorithm 1): one subregion inherits the query’s label, the other the counterfactual’s. This guarantees that a valid decision tree classifier is available at any point during execution. The quality of intermediate classifiers depends on the ordering of the priority queue $\mathcal{Q}$. In practice, preliminary experiments suggested good anytime performance using breadth-first search (BFS), which distributes the exploration evenly. While the design of alternative priority strategies is a promising research direction to enhance TRA’s anytime performance, the choice of exploration order does not impact the total number of queries required for exact reconstruction —this is solely determined by the algorithm’s divide-and-conquer structure.
@@ 4.1 Experimental Setup
Table 1: Summary of the datasets used in our experiments. For each dataset, $m$ is the number of features after pre-processing, encompassing $m_{N}$ numerical, $m_{B}$ binary, $m_{C}$ categorical (before one-hot encoding) and $m_{D}$ discrete (ordinal) ones.
@@
Training the target tree-based models. We train two types of tree-based target models implemented in the scikit-learn library [Pedregosa et al., 2011]: decision trees and random forests. For decision trees, we experiment with varying max_depth parameters ranging from 4 to 10, as well as trees without maximum depth constraint (max_depth set to None). The random forests experiments focus on the COMPAS dataset, employing different numbers of trees (5, 25, 50, 75 and 100) to assess scalability and robustness. To prevent overfitting, we utilize the validation set for hyperparameter tuning and apply cost-complexity pruning where applicable. All the details of training procedures and hyperparameters configurations are discussed in Appendix B.1.
@@
Complete experimental results across all configurations of CF and DualCF are reported in Appendix B.2. They demonstrate that fitting surrogate models of the same hypothesis class facilitates the extraction of both decision trees and random forests. However, knowledge of their hyperparameters does not provide any advantage to either attack. Finally, the non-optimal counterfactuals provided by DiCE lead to better fidelity results than the optimal ones computed by OCEAN for these two attacks. This can be explained by their heuristic nature, which leads to the building of more diverse counterfactuals, not necessarily lying next to a decision boundary. In the next section, we display results only for the best-performing configuration for both CF and DualCF, achieved by training a surrogate model of the same hypothesis class and using DiCE counterfactuals.
Evaluation. We assess each model extraction attack using two metrics: fidelity and number of queries made to the prediction and counterfactual oracle API during the attack. Fidelity measures the proportion of inputs for which the extracted model agrees with the target model, quantifying attack success. We compute fidelity over 3000 points sampled uniformly from the input space, providing a broad evaluation across the entire feature domain. For completeness, we also report fidelity measured on a test set (i.e., drawn from the data distribution) for our experiments using random forest target models, in Appendix B.3, which shows the same performance trends.
Counterfactual oracle. As discussed in Section 3.2, TRA requires the use of a locally optimal counterfactual oracle. Since global optimality is a sufficient condition, we use the popular OCEAN oracle for simplicity in our main experiments with TRA. In Appendix C, we report the performance of TRA using a simple heuristic oracle that produces locally optimal explanations. These experiments show that the choice of oracle (optimal or heuristic) has minimal impact on reconstruction performance.
All experiments are run on a computing cluster with homogeneous nodes using Intel Platinum 8260 Cascade Lake @ 2.4GHz CPU. Each run uses four threads and up to $4$GB of RAM each (multi-threading is only used by the OCEAN oracle). We repeat each experiment five times with different random seeds and report average values. The source code to reproduce all our experiments and figures is accessible at https://github.com/vidalt/Tree-Extractor, under a MIT license.
@@ 4.2 Results
Result 1. TRA outperforms existing approaches in terms of number of queries and anytime fidelity to extract decision trees. Figure 3 presents the average fidelity of surrogate models obtained from the four studied model extraction attacks on decision trees, as a function of the number of queries, for the Adult and COMPAS datasets. Here, each point on a curve is the mean fidelity over all 40 extraction tasks (eight tree depths $\times$ five seeds), so when it reaches 1.00, all target trees were perfectly reconstructed. Results for additional datasets, provided in Figure 11 (Appendix B.4), exhibit the same trends. Across all cases, TRA consistently achieves higher fidelity for any fixed query budget and converges to perfect fidelity orders of magnitude faster. Notably, unlike CF and DualCF, TRA and PathFinding provide formal guarantees of functional equivalence.
Result 2. TRA achieves state-of-the-art performance for functionally equivalent extraction of decision trees. Figure 4 shows the number of queries required by PathFinding and TRA to achieve functionally equivalent model extraction, as a function of the number of nodes in the target models. Results for the Adult and COMPAS datasets are presented, with additional datasets provided in Figure 12 (Appendix B.4). TRA consistently requires orders of magnitude fewer queries than PathFinding to reconstruct the target models with perfect fidelity.
@@
Result 3. TRA theoretically and empirically outperforms existing approaches to extract random forests. Figure 5 presents the average fidelity of the three considered model extraction attacks against random forests, plotted against the number of performed queries for the COMPAS dataset. The results show that TRA achieves higher fidelity with fewer counterfactual queries and converges significantly faster to perfect fidelity. Moreover, TRA is the only attack that certifies functional equivalence for tree ensembles. Additional results in Figure 13 (Appendix B.4) indicate that TRA scales efficiently with the size of the target random forest, as the number of required queries grows sub-linearly with the total number of nodes. This is due to structural redundancies in large forests, allowing the extracted model to be represented with perfect fidelity as a more compact decision tree [Vidal and Schiffer, 2020].
Result 4. TRA is effective with either globally or locally optimal counterfactuals. We report in Appendix C results from experiments where TRA is run with a simple oracle generating heuristic counterfactuals that are only locally optimal. The results show that TRA’s performance is largely unaffected by the lack of global optimality. In some cases, locally optimal counterfactuals even improve early-stage (anytime) performance by introducing greater diversity in the explored input space. These results highlight the practical applicability of TRA in real-world scenarios, where an attacker—given access to any counterfactual oracle—can use simple query access to the target model to iteratively refine counterfactuals via line search, making them locally optimal and suitable to conduct an efficient model extraction attack.

### v3

@@ 1 Introduction
While MLaaS platforms provide accessible and scalable ML solutions, they must address the growing demand for explainability in their decision-making processes. Regulatory frameworks such as the EU AI Act’s Article 13111https://artificialintelligenceact.eu/article/13/ further mandate greater transparency across a wide range of applications. In response, MLaaS providers increasingly incorporate explainability techniques to elucidate model behavior and ensure fairness. Notably, counterfactual explanations specify the changes an input example must undergo to alter its prediction, thereby directly offering a form of recourse in many real-life applications. However, studies have shown that querying a model’s explanations can enable attackers to replicate its parameters and architecture, effectively copying the original model [Tramèr et al., 2016, Wang et al., 2022, Aïvodji et al., 2020, Öksüz et al., 2024]. This reveals a critical tension between the need for transparency and the protection of model integrity and intellectual property.
@@ 2 Online Discovery, Model Extraction Attacks and Competitive Analysis
Competitive Analysis. A central tool for analyzing online discovery problems is competitive analysis [Karlin et al., 1986], which compares the performance of an online algorithm — one that adapts its decisions based solely on information acquired so far — to an optimal offline algorithm with complete foresight. Formally, we measure the ratio between: (i) the complexity (e.g., number of queries, computational cost) incurred by the online algorithm to achieve its goal and (ii) the minimal complexity that an offline algorithm, with complete foresight, would require to accomplish the same task. A constant ratio implies a constant-competitive algorithm; in more complex settings, the ratio may grow with problem parameters.
@@ 3.2 Tree Reconstruction Attack algorithm
We now introduce our proposed Tree Reconstruction Attack (TRA), detailed in Algorithm˜1. TRA is a divide-and-conquer based algorithm that aims to reconstruct a decision tree $f_{n}$ with $n$ split levels by systematically exploring the input space $\mathcal{X}$, using only a black-box API returning predictions and locally optimal counterfactuals. A split level is defined as a particular value for a given feature that divides the input space into two subspaces. Note that multiple nodes within different branches of a decision tree can share the same split level.
Algorithm Overview. TRA operates by maintaining a query list $\mathcal{Q}$ that initially contains the entire input space $\mathcal{X}$. The algorithm iteratively processes each input subset $\mathcal{E}\subseteq\mathcal{X}$ from $\mathcal{Q}$, until $\mathcal{Q}$ is empty, ensuring that all decision boundaries of the target model $f$ are identified and replicated in the reconstructed tree. More precisely, at each iteration, TRA first retrieves the subset $\mathcal{E}$ on top of the priority queue $\mathcal{Q}$. It computes its geometric center $x$ using the centercenter function (line 5). It then queries the oracle $\mathcal{O}_{d}$ with the target model $f$, input $x$, and subset $\mathcal{E}$ to obtain a counterfactual explanation $x^{\prime}=\mathcal{O}_{d}(f,x,\mathcal{E})$ (line 8). The set of feature indices where $x^{\prime}$ differs from $x$, i.e., $\{i\mid x^{\prime}_{i}\neq x_{i}\}$ is consequently identified. For each differing feature $i$, TRA splits the input subset $\mathcal{E}$ into two subspaces based on the split value $x^{\prime}_{i}$ (split function, detailed in Algorithm 2 in the Appendix B). The resulting subspaces are added to $\mathcal{Q}$ for further exploration (line 9). If no counterfactual explanation $x^{\prime}$ exists within $\mathcal{E}$ , TRA assigns the label $y=f(x)$ to $\mathcal{E}$, indicating that it corresponds to a leaf node in the reconstructed tree (line 11).
@@
[…]
@@
9:   $\mathcal{Q}\leftarrow insert(\textsc{split}(\mathcal{E},x,x^{\prime}))$ {Split $\mathcal{E}$ and add the resulting subspaces to $\mathcal{Q}$ (Alg. 2)}
@@
Let $f_{n}$ be a decision tree with $n$ split levels across a $m$-dimensional input space $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{m}$. Denote $s_{i}$ as the number of split levels in $f_{n}$ over the $i$-th feature, such that $\sum_{i=1}^{m}s_{i}=n$. The worst-case complexity of Algorithm˜1 is $O\left(\prod_{\begin{subarray}{c}i=1\end{subarray}}^{m}(s_{i}+1)\right)$.
@@
The worst-case complexity of Algorithm˜1 is $O\left(\left(1+\frac{n}{m}\right)^{m}\right)$.
The proofs of Proposition 3.4 and Corollary 3.5 are provided in the Appendix A. Proposition 3.4 establishes a first simple upper bound on the complexity of Algorithm˜1. Intuitively, consider a two-dimensional decision boundary that partitions the space in a chessboard-like pattern of size $s_{1}\times s_{2}$. A comprehensive mapping of such space necessitates at least $s_{1}\times s_{2}$ queries (one in each sub-square), a requirement that holds for multi-class classification scenarios and in high-dimension.
@@
$C_{TRA}^{(n,m)}=\frac{2\prod_{j=1}^{m}(s_{j}+1)-1}{n+1}\leq\frac{2\left(1+\frac{n}{m}\right)^{m}-1}{n+1},$
@@
The proofs of Propositions 3.6 and 3.7 are provided in Appendix A. Proposition 3.6 provides the competitive ratio achieved by the TRA algorithm, while Proposition 3.7 establishes that the choice of query position does not affect the competitive ratio for any divide-and-conquer algorithm iteratively partitioning the input space. These propositions demonstrate that TRA not only offers a competitive approach to tree reconstruction under various query selection strategies, but also sets a theoretical limit that cannot be surpassed by other methods with a similar divide-and-conquer structure.
Anytime Behavior. Since the query budget is often not known in advance and may vary depending on the target model, it is crucial for an extraction attack to operate in an anytime fashion—that is, to produce a usable classifier even if interrupted before completion. TRA satisfies this property by assigning provisional labels at each input space split (line 9 of Algorithm 1): one subregion inherits the query’s label, the other the counterfactual’s. This guarantees that a valid decision tree classifier is available at any point during execution. The quality of intermediate classifiers depends on the ordering of the priority queue $\mathcal{Q}$. In practice, preliminary experiments suggested good anytime performance using breadth-first search (BFS), which distributes the exploration evenly. While the design of alternative priority strategies is a promising research direction to enhance TRA’s anytime performance, the choice of exploration order does not impact the total number of queries required for exact reconstruction —this is solely determined by the algorithm’s divide-and-conquer structure. Finally, we note that the fraction of total volume corresponding to the leaf regions that have already been fully explored by TRA at a given iteration directly lower bounds the proportion of feature space for which functional equivalence can be guaranteed in an anytime manner. In the case of a uniform data distribution over the feature space, this value also lower bounds the anytime surrogate fidelity, and one could use it to early stop TRA as soon as a target fidelity level is achieved.
@@ 4.1 Experimental Setup
Table 1: Summary of the datasets used in our experiments. For each dataset, $m$ is the number of features after pre-processing, encompassing $m_{N}$ numerical, $m_{B}$ binary, $m_{C}$ categorical (before one-hot encoding) and $m_{D}$ discrete (ordinal) ones. Each of the $m_{C}$ categorical features is one-hot encoded into $c_{j}$ binary dimensions, where $c_{j}$ is the number of categories of feature $j$. As a result, the total number of features becomes: $m=m_{N}+m_{B}+m_{D}+\sum_{j=1}^{m_{C}}c_{j}$.
@@
Training the target tree-based models. We train two types of tree-based target models implemented in the scikit-learn library [Pedregosa et al., 2011]: decision trees and random forests. For decision trees, we experiment with varying max_depth parameters ranging from 4 to 10, as well as trees without maximum depth constraint (max_depth set to None). The random forests experiments focus on the COMPAS dataset, employing different numbers of trees (5, 25, 50, 75 and 100) to assess scalability and robustness. To prevent overfitting, we utilize the validation set for hyperparameter tuning and apply cost-complexity pruning where applicable. All the details of training procedures and hyperparameters configurations are discussed in Appendix C.1.
@@
Complete experimental results across all configurations of CF and DualCF are reported in Appendix C.2. They demonstrate that fitting surrogate models of the same hypothesis class facilitates the extraction of both decision trees and random forests. However, knowledge of their hyperparameters does not provide any advantage to either attack. Finally, the non-optimal counterfactuals provided by DiCE lead to better fidelity results than the optimal ones computed by OCEAN for these two attacks. This can be explained by their heuristic nature, which leads to the building of more diverse counterfactuals, not necessarily lying next to a decision boundary. In the next section, we display results only for the best-performing configuration for both CF and DualCF, achieved by training a surrogate model of the same hypothesis class and using DiCE counterfactuals.
Evaluation. We assess each model extraction attack using two metrics: fidelity and number of queries made to the prediction and counterfactual oracle API during the attack. Fidelity measures the proportion of inputs for which the extracted model agrees with the target model, quantifying attack success. We compute fidelity over 3000 points sampled uniformly from the input space, providing a broad evaluation across the entire feature domain. For completeness, we also report fidelity measured on a test set (i.e., drawn from the data distribution) for our experiments using random forest target models, in Appendix C.3, which shows the same performance trends.
Counterfactual oracle. As discussed in Section 3.2, TRA requires the use of a locally optimal counterfactual oracle. Since global optimality is a sufficient condition, we use the popular OCEAN oracle for simplicity in our main experiments with TRA. In Appendix D, we report the performance of TRA using a simple heuristic oracle that produces locally optimal explanations. These experiments show that the choice of oracle (optimal or heuristic) has minimal impact on reconstruction performance.
All experiments are run on a computing cluster with homogeneous nodes using Intel Platinum 8260 Cascade Lake @ 2.4GHz CPU. Each run uses four threads and up to 44GB of RAM each (multi-threading is only used by the OCEAN oracle). We repeat each experiment five times with different random seeds and report average values. The source code to reproduce all our experiments and figures is accessible at https://github.com/vidalt/Tree-Extractor, under an MIT license.
@@ 4.2 Results
Result 1. TRA outperforms existing approaches in terms of number of queries and anytime fidelity to extract decision trees. Figure 3 presents the average fidelity of surrogate models obtained from the four studied model extraction attacks on decision trees, as a function of the number of queries, for the Adult and COMPAS datasets. Here, each point on a curve is the mean fidelity over all 40 extraction tasks (eight tree depths $\times$ five seeds), so when it reaches 1.00, all target trees were perfectly reconstructed. Results for additional datasets, provided in Figure 12 (Appendix C.4), exhibit the same trends. Across all cases, TRA consistently achieves higher fidelity for any fixed query budget and converges to perfect fidelity orders of magnitude faster. Notably, unlike CF and DualCF, TRA and PathFinding provide formal guarantees of functional equivalence.
Result 2. TRA achieves state-of-the-art performance for functionally equivalent extraction of decision trees. Figure 4 shows the number of queries required by PathFinding and TRA to achieve functionally equivalent model extraction, as a function of the number of nodes in the target models. Results for the Adult and COMPAS datasets are presented, with additional datasets provided in Figure 13 (Appendix C.4). TRA consistently requires orders of magnitude fewer queries than PathFinding to reconstruct the target models with perfect fidelity.
@@
Result 3. TRA theoretically and empirically outperforms existing approaches to extract random forests. Figure 5 presents the average fidelity of the three considered model extraction attacks against random forests, plotted against the number of performed queries for the COMPAS dataset. The results show that TRA achieves higher fidelity with fewer counterfactual queries and converges significantly faster to perfect fidelity. Moreover, TRA is the only attack that certifies functional equivalence for tree ensembles. Additional results in Figure 14 (Appendix C.4) indicate that TRA scales efficiently with the size of the target random forest, as the number of required queries grows sub-linearly with the total number of nodes. This is due to structural redundancies in large forests, allowing the extracted model to be represented with perfect fidelity as a more compact decision tree [Vidal and Schiffer, 2020].
Result 4. TRA is effective with either globally or locally optimal counterfactuals. We report in Appendix D results from experiments where TRA is run with a simple oracle generating heuristic counterfactuals that are only locally optimal. The results (also provided within Figure 6 for the COMPAS dataset) show that TRA’s performance is largely unaffected by the lack of global optimality. In some cases, locally optimal counterfactuals even improve early-stage (anytime) performance by introducing greater diversity in the explored input space. This highlights the practical applicability of TRA in real-world scenarios. Indeed, any valid but possibly sub-optimal counterfactual explanation can be post-processed (through bisection line-searches over each feature using simple prediction queries) into a locally optimal one lying on the decision boundary. Therefore, as long as the counterfactual oracle is reliable (i.e., returns a counterfactual whenever one exists), the returned explanation reveals a direction in which the prediction changes, which is then sufficient to locate a nearby decision boundary and conduct an efficient model extraction attack with TRA.
(a) Anytime performance of TRA: average surrogate fidelity as a function of the number of performed queries.
(b) Performance of TRA to achieve functional equivalence: each point represents the number of queries needed to fully reconstruct the trees.
Figure 6: Comparison of the performance of TRA using either OCEAN or a simpler heuristic counterfactual oracle (Algorithm 3 in Appendix D) to extract decision trees (COMPAS dataset).
Result 5. TRA often requires far fewer queries in practice than its theoretical worst-case bound. Across all datasets we evaluated, the number of counterfactual queries TRA uses to recover a functionally equivalent model is substantially below the worst case of Proposition 3.4. Table 2 reports the empirical query counts and the corresponding theoretical worst-case numbers for depth-9 decision trees (means over multiple random seeds). As evidenced, the former are consistently orders of magnitude smaller than the latter, suggesting that TRA is, on average, considerably more efficient than its theoretical worst-case.
Table 2: Empirical counterfactual query counts used by TRA versus the theoretical worst-case bound from Proposition 3.4 on decision trees of maximum depth 9. Empirical values are averaged over multiple random seeds.
Dataset Empirical # queries Worst-case # queries (Prop. 3.4)
SPerformance $1.16e+03$ $3.02e+08$
Adult $3.70e+04$ $2.45e+14$
German Credit $5.18e+01$ $2.86e+03$
Credit Card $6.97e+04$ $6.49e+11$
COMPAS $1.53e+02$ $1.07e+03$